# An Introduction on Solar Imaging Data Analytic Challenges and Opportunities for Statisticians

### v1

@@ Abstract
We give a gentle introduction to solar imaging data, focusing on challenges and opportunities of data-driven approaches for solar eruptions. The various solar phenomena prediction problems that might benefit from statistical methods are presented. Available data and software will be described. State-of-art solar eruption forecasting with data driven approaches are summarized and discussed. Based on the characteristics of the datasets and state-of-art approaches, we point out several promising directions to explore from statistical modeling and computational perspectives.
@@ 1.1 Introduction to Space Weather and Various Solar Phenomena
Broadly speaking, space weather refers to variable conditions in the solar system environment produced by the Sun’s activity on relatively short time scales (e.g., Baker, 1998; Schwenn, 2006). Long-term effects in space weather on the time scales longer that several solar rotations refer to as space climate (e.g., Mursula et al., 2007). The relation between the solar and geomagnetic activity (solar-terrestrial connections) was noted after Schwabe (1843) discovered the 10-year sunspot cycle, when Sabine (1852), Wolf (1852), Gautier (1852), and Lamont (see, Reslhuber, 1852) reported on sunspot cycle periodicity in variations of Earth geomagnetic field. Large variations in the magnetic field of Earth were noted in relation with the first white-light solar flare, observed by Carrington (1859) and Hodgson (1859). This flare caused one of the most extreme space weather events on record (so called, Carrington event) with wide-spread sightings of aurora borealis, which lasted for about 7 days and was observed as far south as Cuba and Jamaica (see a compillation of several observations in USA, Europe, and Asia, Various, 1859, 1860b, 1860c). For a map of aurora sightings for this event, see Hayakawa et al. (2019). The extended period of Aurora activity started a few days before the Carrington flare, which suggests that multiple eruptive events may have occurred during this period. The resulting geomagnetic activity had an impact on the telegraph, a global communication infrastructure of the time with the variable magnetic field inducing electric currents in the telegraph wires strong enough to make them extremely hot or even spark fires (see a compillation of several reports, Various, 1860a). The appearance of strong electric currents in telegraph wires in conjunction with aurora has been noted before by many observers (e.g., Barlow, 1849; Various, 1860b). For review of early history of space weather impacts including radio communications, navigation, radar detection, and artificial satellites see Pevtsov (2017) and a visualization in Figure 1.
@@
Modern society’s well-being critically depends on global technological systems. This includes global transmission networks of electric power; gas and oil pipelines, the communication and shipping logistics systems, and the global positioning system. Early examples include power outages and blackouts (Boteler et al., 1998; Love et al., 2022, the most cited is the Hydro-Quebec power grid collapse on 13 March 1989, which resulted in the loss of electric power to more than six million customers for nine hours), failure of communication satellites due to radiation damage by solar energetic particles (SEPs) associated with solar flares and coronal mass ejections (e.g., failure of Intelsat’s Galaxy-15 spacecraft in April 2010), electric discharge associated with a satellite-solar wind interaction originated from an elongated coronal hole (Canadian Anik-E1 and E2 communication satellites, Lam et al., 2012), or the enhanced atmospheric drag resulting from a major solar flare/CME event that caused a premature reentry of 38 out of 49 Starlink satellites launched by SpaceX (Berger et al., 2023). In August 1972, series of solar flares and associated geomagnetic storms led to widespread communications disturbances and power outages. These events include a CME with the shortest ever Sun–Earth transit time of just 14.6 hours, an average speed of 2850 km/s and a shock speed at 1 AU 1700 km/s (Zastenker et al., 1978). On 4 August, the geomagnetic storm caused a nearly instantaneous detonation of dozens of sea mines installed by the US during the Vietnam war south of Hai Phong (Knipp et al., 2018). The 4 August flare, which occurred between the Apollo 16 (Apr.16-27, 1972) and Apollo 17 (Dec. 7-11, 1972) missions, led to a significant increase in the radiation level in the interplanetary space. Later estimates (Townsend et al., 1991) indicated that astronauts inside the spacecraft traveling to the Moon would have received a radiation exposure that would exceed annual limits and approach career limits for the skin and ocular lens. The average bone marrow dose equivalent would have exceed the recommended annual limit. For the astronauts outside the spacecraft (extravehicular activity, EVA, or on the moonwalk), the radiation dose equivalents would have been clinically significant including nausea, vomiting, very high probability of cataract formation, a slight increase in the probability of death. Slightly smaller doses were estimated for a flare event in October 1989. For more detail on this topic of impact of SEP radiation on interplanetary travel see Lockwood and Hapgood (2007, and references therein). Finally, in 1967, a large solar flare nearly triggered a nuclear war after it brought down the US early warning radar system with radio noise, causing some military commanders to suspect Soviet jamming in preparation for a nuclear attack (Knipp et al., 2016).
@@
Thus, it should not be surprising that the space weather risk management has become a National priority. The threat-assessment report by the Lloyd’s insurance company (Lloyd’s, 2013) concludes that extreme events could cause $0.6 – 2.6 trillion in damage with a recovery time of months. An earlier report by the National Research Council (Baker et al., 2009) arrived at similar conclusions. More recent studies suggest that in case of an extreme solar and geomagnetic event, daily costs to the world economy would be several billions USD not counting human costs. While extreme space weather events are rare, their damage could be catastrophic. Some studies estimate that the “economic costs associated with these extreme events have been heralded as being as high as $1–2 trillion in the first year, equivalent to a so-called ‘global Hurricane Katrina’ ” (Oughton et al., 2017). Due to the global nature of space weather, the impact will affect a broad swath of the world economy including power distribution, transportation, communications, satellite infrastructure, aviation, and global positioning systems (Eastwood et al., 2017). Besides, penetrating radiation is one of the main obstacles to human exploration of Mars.
@@
While the most dramatic impacts of space weather are due to solar eruptive events, namely flares and CMEs, many other solar phenomena may also play important role. For example, the coronal holes – areas of open magnetic field, could be associated with the recurrent geomagnetic storms of a moderate intensity. The level of the solar UV and EUV radiation, which correlates with solar active regions, heats the Earth’s upper atmosphere, increasing its density and increasing atmospheric drag on the Low Earth Orbit (LEO) satellites. Similarly, the increase in total electron content of the ionosphere affects the speed of radio wave propagation and consequently the accuracy of the global positioning system.
@@
Solar flare (e.g., Bruzek and Durrant, 1977) is a sudden increase in electromagnetic radiation, typically in the range of UV to soft X-rays, from the solar atmosphere (mostly the chromosphere and corona). This radiation is primarily thermal in nature coming from regions of localized heating caused by magnetic energy release. Acceleration of charged particles in flaring regions can also produce non-thermal radiation such as hard X-rays. In the visible light (e.g., H spectral line, 656.3 nm, the flare importance is defined by its relative brightness (F/faint, N/normal, and B/bright) and total area (S/subclass, 1-4 classes). In X-ray, the flares are characterized by the maximum intensity in 0.1-0.8 nm wavelength band:
Class Maximum Intensity Range A (1-9) 10-5 erg/(cm s) B (1-9) 10-4 erg/(cm s) C (1-9) 10-3 erg/(cm s) M (1-9) 10-2 erg/(cm s) X (1-9) 10-1 erg/(cm s) The A-class flares are near the background energy level. A- and B-class flares are normally not geoeffective. C-class flare may have some geomagnetic impact, and M- and X-class flares are the major flares with a major space weather impact. Major flares identified as Importance 3 and 4 in visible light observations, typically correspond to X-ray class M and X. Hayakawa et al. (2020) found that optical flares with an importance of 3 were associated mostly with X-class (66%) and M-class (30%) flares. In 4% cases, strong H flares with C-class flares.
@@
Coronal Mass Ejection (CME, Webb and Howard, 2012) is a large structure of magnetized plasma expelled from the Sun into the interplanetary space. They are associated with strong flares (C, M and X classes) and eruption of the chromospheric filaments (see Figure 2). Their geoeffectiveness depends on their speed (Srivastava and Venkatakrishnan, 2002) and the orientation of magnetic field. Those situated West of solar central meridian, high velocity (fast), and having southward orientation of magnetic field are more geoeffectve, with magnetic field orientation to be the most important parameter. Some CMEs may not have a clearly identifiable source region (so called, stealth CMEs, Ma et al., 2010). They could be associated with the eruption of magnetic structures of filament channels without significant filament material in it (e.g., Pevtsov et al., 2012).
@@
Chromospheric filament (prominance at the limb, Gibson, 2018) is a “cloud” of cooler and denser plasma suspended in the chromosphere and corona by the magnetic field. The vast majority of CMEs are associated with filament channels and early filament rise was found to proceed the CME initiation (see, Berezin et al., 2023, and references therein). Orientation of magnetic field and its handedness (chirality, or sign of magnetic helicity) was found to provide a good representation of southward orientation of magnetic field in CMEs. Filament are located along the polarity inversion line of large-scale magnetic fields at the photosphere.
@@
Coronal hole (CH, Cranmer, 2009) can be identified in coronal images as the darkest (lowest intensity) areas on the solar disk, which can persist for several solar rotations. Historically, CHs are identified using observations in He I 1083.0 nm (near infrared) although their appearance is not as obvious as in extreme UV (EUV) or X-ray images. CHs are associated with weak mostly unipolar magnetic field opened to the interplanetary space. CHs are the source of fast solar wind. Because CHs may exist for several solar rotations, they are associated with the recurrent geomagnetic storms occurring on Earth with the period of solar synodic rotation (about 27.2753 days). The geomagnetic storms associated with CHs are typically moderate in strength. Perhaps, the largest space weather risk from fast-speed streams originating from the coronal holes comes from satellite charging. The resulting electrostatic discharge can damage electronics, phantom commands, cause loss of operations, and in some cases total satellite loss (Horne et al., 2018).
@@
Solar Energetic Particles (SEPs, Reames, 2021). SEPs are the radiation events associated with high-energy particles originating from the Sun in the energy range from about 10 KeV (kilo electron volts) to 10 GeV and lasting from hours to days. Their composition is mostly electrons and protons, but heavy elements (He through Au/Pb) have been identified as well. SEP events are classified as impulsive or gradual. The former originates as the result of magnetic reconnection in solar corona at the location of flares, while the latter are the result of particle acceleration in the CME-driven shocks. Gradual SEPs are also the long-duration events. Some SEPs have sufficient energy to cause localized increase in the radiation at the Earth surface, so called, Ground-Level Events (or Ground-Level Enhancements, GLEs).
@@ 1.2 State-of-Art Solar Events Forecasting with Physics
The state-of-the-art solar event forecasting has been based on physical and empirical models. Empirical solar flare prediction models primarily focuses on parameterizing relationships between the active region’s photospheric magnetic field (e.g., magnetic topology, Lorentz force, free energy, helicity, etc.) and identifying relationships between these parameters and solar flare activity (e.g., McIntosh, 1990; Falconer et al., 2002a; Leka and Barnes, 2003; Schrijver, 2007; Fisher et al., 2012; Moore et al., 2012). The state-of-the-art in physics-based solar flare prediction involves realistic MHD modeling of active region evolution by driving the model’s inner boundary with time-series magnetic field measurements, allowing for the self-consistent modeling of the pre-flare energy build-up process (e.g., Cheung and DeRosa, 2012; Jiang et al., 2016). However, due to the limitations in both physical realism, computational cost and observational data, this advanced modeling remains at the research level and is not yet applicable for operational solar flare prediction.
@@ 1.3 Promises of Data-Driven Approaches
The current space weather forecasting based on physical models is far from reliable: the forecasting window is only minutes and the accuracy is low. Recently, with much more data becoming available, data-driven approaches are gaining attention in the space science community, see Leka and Barnes (2018) and Camporeale (2019) for a review. How to make the best use of the large amount of data available to provide reliable real-time forecasting of space weather events is one of the major questions for scientists in the field. Recent work published in the Space Weather journal (Bobra and Couvidat, 2015; Nishizuka et al., 2018; Chen et al., 2019; Wang et al., 2020; Jiao et al., 2020; Nishizuka et al., 2021) shows that it is highly promising to extend the solar flare forecast time scale from minutes to days using machine learning methods. These results are highly encouraging for the field and provide a benchmark for our future studies. The authors of this paper have published multiple papers on adopting existing ML methods and developing novel statistical methods for solar flare forecasting in the past few years. More specifically, Chen et al. (2019); Wang et al. (2020) adopt the Long Short Term Memory (LSTM) neural network to classify strong solar flare events, demonstrating that constructing precursors of solar flare events from our set of predictors is feasible. In Jiao et al. (2020), we combined the idea of a mixture model and the LSTM to propose a mixed LSTM regression model that predicts the flaring label and flare intensity jointly in one single optimization problem. This results in improved performance in strong flare forecasting. In Sun et al. (2021), we adapt ideas from spatial statistics and topological data analysis to construct physically interpretable predictors of solar flare events, further improving strong flare forecasting performances. In Sun et al. (2022b); Aktukmak et al. (2022), we explore the potential of combining heterogeneous sources of data (data from different instruments from two different solar cycles) for solar flare forecasting, thus seeking to mediate the sample size limitation in rare events (strong solar flare) prediction. In Sun et al. (2023b), we propose a new tensor regression model that combines information from multiple sources of data, accommodating their spatial, temporal, and spectral continuity and sparsity, and propose an efficient computational algorithm to solve the problem.

### v2

# Solar Imaging Data Analytics: A Selective Overview of Challenges and Opportunities

@@ Abstract
We give a gentle introduction to solar imaging data, focusing on the challenges and opportunities of data-driven approaches for solar eruptions. The various solar phenomena prediction problems that might benefit from statistical methods are presented. Available data prodcuts and softwares are described. State-of-the-art solar eruption forecasting models with data-driven approaches are summarized and discussed. Based on the characteristics of the datasets and state-of-the-art approaches, we point out several promising directions to explore from statistical modeling and computational perspectives.
@@ 1.1 Introduction to Space Weather and Various Solar Phenomena
Broadly speaking, space weather refers to variable conditions in the solar system environment produced by the Sun’s activity on relatively short time scales (e.g., Baker, 1998; Schwenn, 2006). Long-term effects in space weather on time scales longer than several solar rotations are referred to as space climate (e.g., Mursula et al., 2007). The relation between the solar and geomagnetic activity (solar-terrestrial connections) was noted after Schwabe (1843) discovered the 10-year sunspot cycle, when Sabine (1852), Wolf (1852), Gautier (1852), and Lamont (see, Reslhuber, 1852) reported on sunspot cycle periodicity in variations of Earth geomagnetic field. Significant variations in the magnetic field of Earth were noted concerning the first white-light solar flare, observed by Carrington (1859) and Hodgson (1859). This flare caused one of the most extreme space weather events on record (the so-called Carrington event). Widespread sightings of aurora borealis, which lasted for about seven days and were observed as far south as Cuba and Jamaica (see a compilation of several observations in USA, Europe, and Asia, Silliman, 1859, 1860b, 1860c). For a map of aurora sightings for this event, see Hayakawa et al. (2019). The extended period of Aurora activity started a few days before the Carrington flare, which suggests that multiple eruptive events may have occurred during this period. The resulting geomagnetic activity had an impact on the telegraph, a global communication infrastructure of the time with the variable magnetic field induces electric currents in the telegraph wires strong enough to make them extremely hot or even spark fires (see a compilation of several reports, Silliman, 1860a). The appearance of strong electric currents in telegraph wires in conjunction with aurora has been noted before by many observers (e.g., Barlow, 1849; Silliman, 1860b). For a review of the early history of space weather impacts, including radio communications, navigation, radar detection, and artificial satellites, see Pevtsov (2017) and visualization in Figure 1.
@@
Modern society’s well-being critically depends on global technological systems. This includes global transmission networks of electric power, gas and oil pipelines, communication and shipping logistics systems, and the global positioning system. Early examples include power outages and blackouts (Boteler et al., 1998; Love et al., 2022, the most cited is the Hydro-Quebec power grid collapse on 13 March 1989, which resulted in the loss of electric power to more than six million customers for nine hours), failure of communication satellites due to radiation damage by solar energetic particles (SEPs) associated with solar flares and coronal mass ejections (e.g., failure of Intelsat’s Galaxy-15 spacecraft in April 2010), electric discharge associated with a satellite-solar wind interaction originated from an elongated coronal hole (Canadian Anik-E1 and E2 communication satellites, Lam et al., 2012), or the enhanced atmospheric drag resulting from a major solar flare or coronal mass ejection (CME) event that caused a premature reentry of 38 out of 49 Starlink satellites launched by SpaceX (Berger et al., 2023). In August 1972, a series of solar flares and associated geomagnetic storms led to widespread communications disturbances and power outages. These events include a CME with the shortest ever Sun–Earth transit time of just 14.6 hours, an average speed of 2850 km/s and a shock speed at 1 AU 1700 km/s (Zastenker et al., 1978). On 4 August, the geomagnetic storm caused a nearly instantaneous detonation of dozens of sea mines installed by the US during the Vietnam War south of Hai Phong (Knipp et al., 2018). The 4 August flare, which occurred between the Apollo 16 (Apr.16-27, 1972) and Apollo 17 (Dec. 7-11, 1972) missions, led to a significant increase in the radiation level in the interplanetary space. Later estimates (Townsend et al., 1991) indicated that astronauts inside the spacecraft traveling to the Moon would have received radiation exposure that would exceed annual limits and approach career limits for the skin and ocular lens. The average bone marrow dose equivalent would have exceeded the recommended annual limit. For the astronauts outside the spacecraft (extravehicular activity, EVA, or on the moonwalk), the radiation dose equivalents would have been clinically significant, including nausea, vomiting, a very high probability of cataract formation, a slight increase in the probability of death. Slightly smaller doses were estimated for a flare event in October 1989. For more detail on this topic of impact of SEP radiation on interplanetary travel see Lockwood and Hapgood (2007, and references therein). Finally, in 1967, a large solar flare nearly triggered a nuclear war after it brought down the US early warning radar system with radio noise, causing some military commanders to suspect Soviet jamming in preparation for a nuclear attack (Knipp et al., 2016).
@@
Thus, it should not be surprising that space weather risk management has become a National priority. The threat-assessment report by the Lloyd’s Insurance company (Lloyd’s, 2013) concludes that extreme events could cause $0.6 – 2.6 trillion in damage with a recovery time of months. An earlier report by the National Research Council (Baker et al., 2009) arrived at similar conclusions. More recent studies suggest that in case of an extreme solar and geomagnetic event, daily costs to the world economy would be several billion USD, not counting human costs. While extreme space weather events are rare, their damage could be catastrophic. Some studies estimate that the “economic costs associated with these extreme events have been heralded as being as high as $1–2 trillion in the first year, equivalent to a so-called ‘global Hurricane Katrina’ ” (Oughton et al., 2017). Due to the global nature of space weather, the impact will affect a broad swath of the world economy, including power distribution, transportation, communications, satellite infrastructure, aviation, and global positioning systems (Eastwood et al., 2017). Besides, penetrating radiation is one of the main obstacles to human exploration of Mars.
@@
While the most dramatic impacts of space weather are due to solar eruptive events, namely flares and CMEs, many other solar phenomena may also play an important role. For example, the coronal holes – areas of open magnetic field, could be associated with recurrent geomagnetic storms of moderate intensity. The level of the solar UV and EUV radiation, which correlates with solar active regions, heats the Earth’s upper atmosphere, increasing its density and increasing atmospheric drag on the Low Earth Orbit (LEO) satellites. Similarly, the increase in total electron content of the ionosphere affects the speed of radio wave propagation and, consequently, the accuracy of the global positioning system.
@@
Solar flare (e.g., Bruzek and Durrant, 1977) is a sudden increase in electromagnetic radiation, typically in the range of UV to soft X-rays, from the solar atmosphere (mostly the chromosphere and corona). This radiation is primarily thermal in nature, coming from regions of localized heating caused by magnetic energy release. Acceleration of charged particles in flaring regions can also produce non-thermal radiation, such as hard X-rays. In the visible light (e.g., H spectral line, 656.3 nm, the flare importance is defined by its relative brightness (F/faint, N/normal, and B/bright) and total area (S/subclass, 1-4 classes). In X-ray, the flares are characterized by the maximum intensity in the 0.1-0.8 nm wavelength band:
Class Maximum Intensity Range A (1-9) 10-5 erg/(cm s) B (1-9) 10-4 erg/(cm s) C (1-9) 10-3 erg/(cm s) M (1-9) 10-2 erg/(cm s) X (1-9) 10-1 erg/(cm s) The A-class flares are near the background energy level. A- and B-class flares are typically not geoeffective. C-class flares may have some geomagnetic impact, and M- and X-class flares are the major flares with a major space weather impact. Major flares identified as Importance 3 and 4 in visible light observations typically correspond to X-ray class M and X. Hayakawa et al. (2020) find that optical flares with an importance of 3 were associated mainly with X-class (66%) and M-class (30%) flares. In 4% cases, strong H flares with C-class flares.
@@
Coronal Mass Ejection (CME, Webb and Howard, 2012) is a large structure of magnetized plasma expelled from the Sun into the interplanetary space. They are associated with strong flares (C, M, and X classes) and eruption of the chromospheric filaments (see Figure 2). Their geoeffectiveness depends on their speed (Srivastava and Venkatakrishnan, 2002) and the orientation of the magnetic field. Those situated west of the solar central meridian, with high velocity (fast) and a southward orientation of the magnetic field, are more geoeffective, with magnetic field orientation being the most important parameter. Some CMEs may not have a clearly identifiable source region (so called, stealth CMEs, Ma et al., 2010). They could be associated with the eruption of magnetic structures of filament channels without significant filament material in it (e.g., Pevtsov et al., 2012).
@@
Chromospheric filament (prominance at the limb, Gibson, 2018) is a “cloud” of cooler and denser plasma suspended in the chromosphere and corona by the magnetic field. The vast majority of CMEs are associated with filament channels, and early filament rise was found to proceed the CME initiation (see, Berezin et al., 2023, and references therein). The orientation of the magnetic field and its handedness (chirality, or sign of magnetic helicity) are found to provide a good representation of the southward orientation of the magnetic field in CMEs. Filaments are located along the polarity inversion line of large-scale magnetic fields at the photosphere.
@@
Coronal hole (CH, Cranmer, 2009) can be identified in coronal images as the darkest (lowest intensity) areas on the solar disk, which can persist for several solar rotations. Historically, CHs have been identified using observations in He I 1083.0 nm (near infrared), although their appearance is not as evident as in extreme UV (EUV) or X-ray images. CHs are associated with weak, mostly unipolar magnetic fields opened to the interplanetary space. CHs are the source of fast solar wind. Because CHs may exist for several solar rotations, they are associated with the recurrent geomagnetic storms occurring on Earth with the period of solar synodic rotation (about 27.2753 days). The geomagnetic storms associated with CHs are typically moderate in strength. Perhaps the largest space weather risk from fast-speed streams originating from the coronal holes comes from satellite charging. The resulting electrostatic discharge can damage electronics, phantom commands, cause loss of operations, and in some cases, total satellite loss (Horne et al., 2018).
@@
Solar Energetic Particles (SEPs, Reames, 2021). SEPs are the radiation events associated with high-energy particles originating from the Sun in the energy range from about 10 KeV (kilo electron volts) to 10 GeV, lasting from hours to days. Their composition is mostly electrons and protons, but heavy elements (He through Au/Pb) have been identified as well. SEP events are classified as impulsive or gradual. The former originates as the result of magnetic reconnection in the solar corona at the location of flares, while the latter is the result of particle acceleration in the CME-driven shocks. Gradual SEPs are also long-duration events. Some SEPs have sufficient energy to cause localized increases in the radiation at the Earth’s surface, so-called Ground-Level Events (or Ground-Level Enhancements, GLEs).
@@ 1.2 State-of-Art Solar Events Forecasting with Physics
State-of-the-art solar event forecasting has been based on physical and empirical models. Empirical solar flare prediction models primarily focus on parameterizing relationships between the active region’s photospheric magnetic field (e.g., magnetic topology, Lorentz force, free energy, helicity, etc.) and identifying relationships between these parameters and solar flare activity (e.g., McIntosh, 1990; Falconer et al., 2002a; Leka and Barnes, 2003; Schrijver, 2007; Fisher et al., 2012; Moore et al., 2012). The state-of-the-art in physics-based solar flare prediction involves realistic MHD modeling of active region evolution by driving the model’s inner boundary with time-series magnetic field measurements, allowing for the self-consistent modeling of the pre-flare energy build-up process (e.g., Cheung and DeRosa, 2012; Jiang et al., 2016). However, due to the limitations in both physical realism, computational cost, and observational data, this advanced modeling remains at the research level. It is not yet applicable for operational solar flare prediction.
@@ 1.3 Promises of Data-Driven Approaches
The current space weather forecasting based on physical models is far from reliable: the forecasting window is only minutes, and the accuracy is low. Recently, with much more data becoming available, data-driven approaches are gaining attention in the space science community; see Leka and Barnes (2018) and Camporeale (2019) for a review. How to make the best use of the large amount of data available to provide reliable real-time forecasting of space weather events is one of the major questions for scientists in the field. Recent work published in the Space Weather journal (Bobra and Couvidat, 2015; Nishizuka et al., 2018; Chen et al., 2019; Wang et al., 2020; Jiao et al., 2020; Nishizuka et al., 2021) shows that it is highly promising to extend the solar flare forecast time scale from minutes to days using machine learning methods. These results are highly encouraging for the field and provide a benchmark for future studies. The authors of this paper have published multiple papers on adopting existing ML methods and developing novel statistical methods for solar flare forecasting in the past few years. More specifically, Chen et al. (2019); Wang et al. (2020) adopt the Long Short Term Memory (LSTM) neural network to classify strong solar flare events, demonstrating that constructing precursors of solar flare events from our set of predictors is feasible. In Jiao et al. (2020), we combined the idea of a mixture model and the LSTM to propose a mixed LSTM regression model that predicts the flaring label and flare intensity jointly in one single optimization problem. This results in improved performance in strong flare forecasting. In Sun et al. (2021), we adapt ideas from spatial statistics and topological data analysis to construct physically interpretable predictors of solar flare events, further improving strong flare forecasting performances. In Sun et al. (2022b); Aktukmak et al. (2022), we explore the potential of combining heterogeneous sources of data (data from different instruments from two different solar cycles) for solar flare forecasting, thus seeking to mediate the sample size limitation in rare events (strong solar flare) prediction. In Sun et al. (2023b), we propose a new tensor regression model that combines information from multiple sources of data, accommodating their spatial, temporal, and spectral continuity and sparsity, and propose an efficient computational algorithm to solve the problem.